Listed below are the top 10 annual salaries​ (in millions of​ dollars) of TV personalities. Find the​ range, variance, and standard deviation for the sample data. Given that these are the top 10​ salaries, do we know anything about the variation of salaries of TV personalities in​ general? 38 37 36 30 19 17 14 13 12.6 12.0



1
Expert's answer
2021-09-30T00:55:30-0400

The range is the difference between the smallest and the highest value.


"38-12.0=26"

The mean is: 

"mean=\bar{x}=\frac{1}{10}(38 +37+ 36+ 30+ 19+ 17""+14 +13+ 12.6+ 12.0)=22.86"

The variance is:

"Variance=s^2=\frac{1}{10-1}((38-22.86)^2+(37-22.86)^2"




"+(36-22.86)^2+(30-22.86)^2+(19-22.86)^2"




"+(17-22.86)^2+(14-22.86)^2+(13-22.86)^2"

"+(12.6-22.86)^2+(12.0-22.86)^2)=122.329333"

The standard variation is


"s=\sqrt{s^2}=\sqrt{122.329333}=11.060259"




"range=26"

"mean=22.86"

"Variance=122.33"

"standard\ deviation=11.06"

No, because the sample is not representative of the whole population.